OpenStudy (anonymous):

35.752 is 43.6% of what number?

hero (hero):

s/of = %/100 35.752/x = 43.6/100 43.6x = 3575.2 x = 3575.2/43.6 = 82

OpenStudy (anonymous):

thanks so much!!

hero (hero):

Oh, I didn't realize you were still here